Q2/please I posted before 2 days but the answer are wrong Please answer question

Pharoah Limited purchased a machine on account on April 2, 2018, at an invoice price of $356,620. On April 4, it paid $2,130 for delivery of the machine. A one-year, $3,970 insurance policy on the machine was purchased on April 5. On April 18, Pharoah paid $7,590 for installation and testing of the machine. The machine was ready for use on April 30.

Pharoah estimates the machine’s useful life will be five years or 6,212 units with a residual value of $73,690. Assume the machine produces the following numbers of units each year: 896 units in 2018; 1,400 units in 2019; 1,405 units in 2020; 1,396 units in 2021; and 1,115 units in 2022. Pharoah has a December 31 year end.

A/ Determine the cost of the machine.

Cost of the machine

$

B/

Calculate the annual depreciation and total depreciation over the asset’s life using: (Round the depreciation cost per unit to the nearest cent. Round answers to 0 decimal places, e.g. 5,275.)

(1) Straight-line method

Year

Depreciable Cost

Depreciation Expense

Accumulated Depreciation

Carrying Amount

2018

$

$

$

$

2019

  

  

  

  

2020

  

  

  

  

2021

  

  

  

  

2022

  

  

  

  

2023

  

  

  

  

(2) Double-diminishing-balance method

Year

Opening Carrying Amount

Depreciation Expense

Accumulated Depreciation

Carrying Amount

2018

$

$

$

$

2019

  

  

  

  

2020

  

  

  

  

2021

  

  

  

  

2022

  

  

  

  

2023

  

  

  

  


(3) Units-of-production method

Year

Units-of-production

Depreciation Expense

Accumulated Depreciation

Carrying Amount

2018

$

$

$

2019

  

  

  

2020

  

  

  

2021

  

  

  

2022

  

  

  


(4) Which method causes net income to be lower in the early years of the asset’s life?

Straight-Line MethodDouble-Diminishing-Balance MethodUnits-of-Production MethodStraight-Line MethodDouble-Diminishing-Balance MethodUnits-of-Production Method
Answer #1

A) Cost of Machine is calculated as follows:- (Amounts in $)

Purchase Cost of machine (Invoice price) 356,620
Delivery charges for machine 2,130
Installation and testing charges 7,590
Total cost of machine 366,340

Therefore total cost of machine is $366,340.

(Insurance expense will not capitalized to the cost of machine).

B) 1) Straight Line Method:-

Depreciation per year = (Cost - Residual Value)/Useful Life in yrs

= ($366,340 - $73,690)/5 yrs = $58,530 per year

Schedule showing annual depreciation (Amounts in $)

Year

Depreciable Cost

Depreciation Expense

Accumulated Depreciation

Carrying Amount

2018

292,650 (366,340-73,690)

39,020 (58,530*8/12)

39,020

327,320 (366,340-39,020)

2019

292,650 (366,340-73,690)

58,530

97,550

268,790 (327,320-58,530)  

2020

292,650 (366,340-73,690)

58,530

156,080

210,260 (268,790-58,530)

2021

292,650 (366,340-73,690)

58,530

214,610

151,730 (210,260-58,530)

2022

292,650 (366,340-73,690)

58,530  

273,140

93,200 (151,730-58,530)

2023

292,650 (366,340-73,690)

19,510 (58,530*4/12)

292,650

73,690 (93,200-19,510)

2) Double-Declining Balance Method:-

Depreciation rate under double declining balance = (1/Useful life in yrs)*2*100

= (1/5 yrs)*2*100 = 40%

Schedule showing annual depreciation (Amounts in $)

Year

Opening Carrying Amount (A)

Depreciation Expense (B = A*40%)

Accumulated Depreciation

Carrying Amount (A-B)

2018

366,340 97,691 (366,340*40%*8/12) 97,691 268,649

2019

268,649

107,460

205,151  

161,189

2020

161,189

64,476  

269,627

96,713  

2021

96,713

38,685

308,312

58,028

2022

58,028

23,211

331,523  

34,817  

2023

34,817

4,642 (34,817*40%*4/12)

336,165

30,175  

3) Units of Production Method:-

Depreciation expense per unit = (Cost - Residual Value)/Useful life in units

= ($366,340 - $73,690)/6,212 units = $47.11 per unit

Schedule showing annual depreciation (Amounts in $)

Year

Units-of-production (A)

Depreciation Expense (A*47.11)

Accumulated Depreciation (B)

Carrying Amount (366,340-B)

2018

896 units 42,211 42,211 324,129

2019

1,400 units

65,954

108,165

258,175

2020

1,405 units

66,190

174,355  

191,985  

2021

1,396 units

65,766

240,121

126,219  

2022

1,115 units

52,528

292,649  

73,691

4) The net income will be lower under that method in which depreciation cost is higher (i.e. double declining balance method). Therefore Double-Diminishing-Balance Method causes net income to be lower in the early years of the asset’s life.